# Quantum thermocouples: nonlocal conversion and control of heat in semiconductor nanostructures

### v1

@@ III.1 Inelastic junctions
FIG. 2: (a) Inelastic-scattering junction, with the region in contact with the heat source sandwiched between two scattering regions ${\cal S}_{\alpha}$, $\alpha=1,2$, which connect it with the conductor terminals $L$ and $R$. (b) Resonant tunneling barriers introduce Lorentzian-shaped filters ${\cal T}_{\alpha}(E)$ that filter the transported electrons. Resonances can be tuned with gate voltages $V_{g\alpha}$. Those below (over) the electrochemical potential induce transport from hot to cold (from cold to hot) regions. (c) A finite power $P$ is generated in asymmetric configurations by generating a current through a load resistor closing the circuit. Here $R_{load}\gg R_{K}$, with resonances of width $\Gamma_{1}=0.2\,\mathrm{meV}$, $\Gamma_{1}=0.1\,\mathrm{meV}$, $T=175\,\mathrm{mK}$, $\Delta T=5\,\mathrm{mK}$.
@@
$I_{L}=\frac{G_{LH}L_{RH}-G_{RH}L_{LH}}{G_{LH}+G_{RH}}\Delta T,$ (13)
with a developed potential $\mu_{H}-\mu=(L_{LH}+L_{RH})\Delta T_{H}/(G_{LH}+G_{RH})$  [^2] The two junctions act as two parallel resistors to the particle current flowing from the hot terminal generated by them.. As shown in Fig. 1(b), $L_{LH}$ and $L_{RH}$ of opposite signs can be achieved for resonant barrier scatterers by tuning gate voltages such that they can both contribute positively to the current in Eq. (13).
@@ III.1.1 Resonant tunneling energy harvesters
The quantum dot energies, $\varepsilon_{l}$, can be tuned with gate voltages, one of them to filter particles below the electrochemical potential (hence contributing to transport into the hot cavity), the other one over it (therefore extracting particles from the cavity) [115, 136], which results in a direct current that can power a load resistor, as shown in Fig. 2(c). The generated current can be scaled up by parallelization in a layered structured of self-assembled quantum dots [115]. Alternatively, quantum wells [133] or superlattices [135] enhance the available currents while keeping the filtering properties (the maximum power is $P\sim 0.1-0.3\,\mathrm{W/cm^{2}}$ at an efficiency around $0.1-0.2\eta_{C}$ at room temperature [135]). The band quantization of one-dimensional wires [206] and two-dimensional heterostrusctures [207] can be used as low-pass filters as mesoscopic versions of thermionic devices, which however compromise the efficiency. Multiple resonance barriers [139] have also been considered which may approach the optimal configuration, with boxcar-shaped transmissions, in terms of efficiency at finite power [203]. The three-terminal composition can also be exploited to enhance the engine performance by allowing for cooperative effects due to having two temperature differences ($T_{L}\neq T_{R}\neq T_{H}$[134].
@@ III.2 Phase coherent thermocouples
To show this, it is useful to partition the junction in three regions: two scattering elements, $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, mediating the transport between the junction and the isothermal terminals $L$ and $R$, and a central region where the conductor couples to the heat source, which is injected coherently into the scattering region. A possible realization is sketched in Fig. 3(a), where the heat source is connected to a scanning probe tip allowing for the local injection of heat at a chosen position and for choosing the coupling strength (and even be disconnected from the conductor). The electron waves scattered into the channels connecting two consecutive scattering elements accumulate a phase that depends on the distances $x$ and $d$ which, after undergoing multiple reflections between them, gives rise to interference processes. This phase can be furthermore controlled: kinetic phases, which are proportional to $k(E)=\sqrt{2m(E-U)}/\hbar$, are tuned with plunger gates acting on the base electrostatic potential, $U$; magnetic phases in annular configurations are tuned via the magnetic flux piercing the ring, $\Phi$ [211, 212].
@@
Equation (13) shows that inelastic thermocouples need to be formed of elements with a finite Seebeck response. A remarkable difference in phase coherent junctions is the possibility to use the hot terminal to induce an extrinsic thermoelectric effect in a conductor that would exhibit no thermoelectricity without the probe [149]. A simple example is a one channel conductor with an energy-independent scattering matrix ${\cal S}_{1}$ connecting terminals $L$ and $R$ to which a hot scanning tip is approached. This is the configuration of Fig. 3(a) with ${\cal S}_{2}=\mathbb{1}$ and ${\cal S}_{1}$ depending on a constant transmission probability. The kinetic phase accumulated in multiple reflections between the barrier and the tip (separated by a distance $x$) introduces an energy dependent $2k(E)|x|$-periodic interference pattern in the three-terminal scattering matrix (composed of the barrier and the tip) sufficient to break the electron-hole symmetry [213]. Meanwhile, the inversion symmetry is broken geometrically by the position of the tip with respect to the barrier. The interference pattern can be further tuned by modifying the position of the tip. This results in a tunable thermoelectric effect as shown in Fig. 3(b), where the current, $I=I_{L}=-I_{R}$, exhibits a series of oscillations with $x$. A similar non-local oscillating behavior was found in tight-binding simulations of a quantum point contact acting as the isothermal conductor [148]. The tunability of the transport properties of these systems allow furthermore for highly efficient thermoelectric rectification [148, 149]. Unfortunately, the heat to work conversion performance is low in these devices, as a consequence of a considerable fraction of the heat being lost to the terminal without thermoelectric scatterer and the transmission probability being sub-optimal.
@@
The performance can however be improved by considering scattering regions with more useful properties ${\cal S}_{\alpha}(E)$ and exploring the capability to engineer the global transmission probabilities of a quantum thermocouple using the electron phase. A system comprised by quantum dots (or double quantum dots) as scattering regions $1$ and $2$, has been proposed [104] as a coherent analog of the inelastic thermocouples [115, 63] discussed in Sec. III.1.1. While this system allows for direct transport between terminals $L$ and $R$, which is in principle detrimental for the thermocouple operation (as discussed in Sec. III), optimizing the different available parameters one finds regimes for which the tip-mediated interference within the conductor becomes destructive and ${\cal T}_{LR}(E)\approx 0$, see Fig. 3(c). Furthermore, this interference can improve the transmission properties of the two junctions giving rise to an enhancement of all performance quantifiers with respect to the inelastic case. In particular, for double quantum dots, the resulting transmissions develop a boxcar-shape [203], attaining efficiencies of $\eta>0.4\eta_{C}$ with high extracted powers (around $70\%$ of the theoretical maximum) and reduced signal to noise ratio [104]. A slight improvement of the performance was also found in tight-binding coherent models of the quantum dot thermocouple [137], which can again be attributed to suppressed transport between isothermal terminals and the favorable modification of the $H-L$ and $H-R$ transmissions.
@@
The manipulation of the electron phase via the Aharonov-Bohm effect [214] in three-terminal coherent ring configurations to design quantum thermocouples has also been explored [130, 215, 150]. The properties of these devices are easily tunable via the magnetic phase $\phi=2\pi\Phi/\Phi_{0}$ accumulated upon performing a loop around the ring, with $\Phi$ being the magnetic flux piercing the ring and $\Phi_{0}=h/e$ the flux quantum [13]. In these setups the electron-hole symmetry is broken by the kinetic phase accumulated between the junctions connecting the ring to the different terminals [216, 217], with broken time-reversal symmetry introducing the mirror asymmetry (via the transmission probabilities ${\cal T}_{ij}(B)\neq{\cal T}_{ji}(B)$) needed for the thermocouple effect, even if the geometric configuration of the system is symmetric and energy-independent [150]. To see this, consider a symmetric ring with all three terminals equally separated by one third of the ring perimeter, so we have ${\cal T}_{\circlearrowright}(\phi)={\cal T}_{12}(\Phi)={\cal T}_{23}(\Phi)={\cal T}_{31}(\Phi)$. A Sommerfeld expansion [218] of the linear response coefficients (valid at low temperatures, $k_{\text{B}}T\ll\mu$ provided the transmission probabilities are smooth around $E=\mu$ [149]) results in a zero-bias current:
$I_{1}=2q_{H}\frac{{\cal T}_{\circlearrowright}(\mu,\phi){\cal T}^{\prime}_{\circlearrowright}(\mu,{-}\phi){-}{\cal T}_{\circlearrowright}(\mu,{-}\phi){\cal T}^{\prime}_{\circlearrowright}(\mu,\phi)}{{\cal T}_{\circlearrowright}(\mu,\phi)+{\cal T}_{\circlearrowright}(\mu,-\phi)}\Delta T_{H},$ (15)
@@
The relaxation of electrons within the conductor can also be modeled via probe terminals by imposing the boundary conditions $I_{p}=J_{p}=0$: the probe adapts its temperature and electrochemical potential, thus behaving as a thermometer, and enable the definition of an effective temperature in different areas of the conductor [220, 221], which also shows an interference pattern [149]. The measurement of local dissipation in one dimensional conductors via scanning probe tips has also been described with more involved models combining the scattering approach with a nonlinear Boltzmann equation for the electrons [222].
III.2.1 Edge states
@@
A way to control which electrons get in contact with the heat source is to use the chirality of quantum Hall edge states [225, 226] in the presence of strong magnetic fields. The quantum Hall regime is indeed a multiterminal effect, and thermoelectric measurements have elegantly demonstrated the chiral propagation of electrons via the energy dissipated in the terminals along the way from the heat source [68, 69].
In addition to the separation of heat and particle flows discussed above, the quantum Hall regime achieves the separation of left moving and right moving electrons on opposite sides of the conductor, making it possible to inject heat only on electrons moving in a particular direction [227, 140, 141, 142, 228], see Fig. 4(a). In the absence of interedge scattering, heat flows only along the edge connected to the hot terminal. Along the other edge, the particle current vanishes. In order to generate a finite thermoelectric response, the two edges need to be coupled by an energy-dependent scattering region, see Fig. 4(b). This can be a gate tunable junction like, e.g., a quantum point contact or a quantum dot, or induced by the difference of kinetic phases in a Mach-Zehnder interferometer [143]. Interestingly, if the scatterer is placed only between the heat source and one of the conductor terminals, as shown in Fig. 4(b), the Seebeck and Peltier effects (which are linked by Onsager reciprocity relations in the absence of a magnetic field) are also separated, i.e., one is finite with the other one vanishing [141]: $\Delta T$ generates a current between $L$ and $R$ (hot particles from $H$ are filtered at the scatterer), but a difference in $\mu_{L}-\mu_{R}$ does not generate a thermoelectric heat current in terminal $H$ (particles injected from $L$ never reach $H$). This effect was related to a possible enhancement of the efficiency at maximum power [165], however it was later shown that the balance established by the not scattered channel restores the usual bounds [203, 147].
The same discussion can be applied to helical states in topological insulators, which preserve time-reversal symmetry but separate electrons with opposite spin polarization. In that case, the injected heat is converted into a pure spin current [141, 144].
@@ III.2.2 Nonequilibrium states
Proposals to observe such effect include the use of quantum Hall edge states that facilitate the preparation of a nonequilibrium distribution: two terminals at different local states (e.g., $f_{C}(E)$ and $f_{H}(E)$) can be connected by a scatterer which mixes their distribution. Electrons in one of the outgoing channels from this scattering region will have a distribution $f_{neq}(E)={\cal T}_{C}(E)f_{C}(E)+{\cal T}_{H}(E)f_{H}(E)$, where ${\cal T}_{i}$ is the transmission probability for electrons injected from terminal $i$ [145, 146]. Electrons in the edge state with the resulting distribution can then be injected into the conductor inducing a finite power if they are scattered at a junction with a thermoelectric response, similarly as the discussion in Sec. III.2.1, see Fig. 4(c). In this model, the fulfillment of the second law can be related to the dissipative currents within the demon, $J_{H}=-J_{C}$, with $\dot{S}_{d}=\dot{S}_{C}+\dot{S}_{H}=J_{H}(T_{C}^{-1}-T_{H}^{-1})\geq-\dot{S}_{s}$. The nonequilibrium situation can also be generated by time-dependent pulses [232]. The addition of channels coupled to different phonon baths have also been discussed to have a similar effect that however cannot be related to a single nonequilibrium state [158].
In experimental configurations, the nonequilibrium situation can be a result of different processes acting on different time scales. This is precisely the regime of operation of hot-carrier solar cells, which can be modelled as a multiterminal conductor [187]. Mesoscopic conductors are then of help for understanding how to extract useful power of nonequilibrium states before they relax in practical setups.

### v2

@@ III.1 Inelastic junctions
FIG. 2: (a) Inelastic-scattering junction, with the region in contact with the heat source sandwiched between two scattering regions ${\cal S}_{\alpha}$, $\alpha=1,2$, which connect it with the conductor terminals $L$ and $R$. (b) Resonant tunneling barriers introduce Lorentzian-shaped filters ${\cal T}_{\alpha}(E)$ that filter the transported electrons. Resonances can be tuned with gate voltages $V_{g\alpha}$. Those below (over) the electrochemical potential induce transport from hot to cold (from cold to hot) regions. (c) A finite power $P$ is generated in asymmetric configurations by generating a current through a load resistor closing the circuit. Here $R_{load}\gg R_{K}$, with resonances of width $\Gamma_{1}=0.2\,\mathrm{meV}$, $\Gamma_{1}=0.1\,\mathrm{meV}$, $T=175\,\mathrm{mK}$, $\Delta T=5\,\mathrm{mK}$, and being $R_{K}=h/e^{2}\approx 25.813\,\mathrm{k\Omega}$ the von Klitzing resistance.
@@
$I_{L}=\frac{G_{RH}L_{LH}-G_{LH}L_{RH}}{G_{LH}+G_{RH}}\Delta T,$ (13)
with a developed potential $\mu_{H}-\mu=-(L_{LH}+L_{RH})\Delta T_{H}/(G_{LH}+G_{RH})$. [^2] The two junctions act as two parallel resistors to the particle current flowing from the hot terminal generated by them. As shown in Fig. 1(b), $L_{LH}$ and $L_{RH}$ of opposite signs can be achieved for resonant barrier scatterers by tuning gate voltages such that they can both contribute positively to the current in Eq. (13).
@@ III.1.1 Resonant tunneling energy harvesters
The quantum dot energies, $\varepsilon_{l}$, can be tuned with gate voltages, one of them to filter particles below the electrochemical potential (hence contributing to transport into the hot cavity), the other one over it (therefore extracting particles from the cavity) [126, 149], which results in a direct current that can power a load resistor, as shown in Fig. 2(c). The generated current can be scaled up by parallelization in a layered structured of self-assembled quantum dots [126]. Alternatively, quantum wells [146] or superlattices [148] enhance the available currents while keeping the filtering properties (the maximum power is $P\sim 0.1-0.3\,\mathrm{W/cm^{2}}$ at an efficiency around $0.1-0.2\eta_{C}$ at room temperature [148]). The band quantization of one-dimensional wires [235] and two-dimensional heterostrusctures [236] can be used as low-pass filters as mesoscopic versions of thermionic devices, which however compromise the efficiency. Multiple resonance barriers [152] have also been considered which may approach the optimal configuration, with boxcar-shaped transmissions, in terms of efficiency at finite power [230]. The three-terminal composition can also be exploited to enhance the engine performance by allowing for cooperative effects due to having two temperature differences ($T_{L}\neq T_{R}\neq T_{H}$) [147].
@@ III.2 Phase coherent thermocouples
To show this, it is useful to partition the junction in three regions: two scattering elements, $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, mediating the transport between the junction and the isothermal terminals $L$ and $R$, and a central region where the conductor couples to the heat source, which is injected coherently into the scattering region. A possible realization is sketched in Fig. 3(a), where the heat source is connected to a scanning probe tip allowing for the local injection of heat at a chosen position and for choosing the coupling strength (and even be disconnected from the conductor). The electron waves scattered into the channels connecting two consecutive scattering elements accumulate a phase that depends on the distances $x$ and $d$ which, after undergoing multiple reflections between them, gives rise to interference processes. This phase can be furthermore controlled: kinetic phases, which are proportional to $k(E)=\sqrt{2m(E-U_{0})}/\hbar$, are tuned with plunger gates acting on the base electrostatic potential, $U_{0}$; magnetic phases in annular configurations are tuned via the magnetic flux piercing the ring, $\Phi$ [240, 241].
@@
Equation (13) shows that inelastic thermocouples need to be formed of elements with a finite Seebeck response. A remarkable difference in phase coherent junctions is the possibility to use the hot terminal to induce an extrinsic thermoelectric effect in a conductor that would exhibit no thermoelectricity without the probe [163]. A simple example is a one channel conductor with an energy-independent scattering matrix ${\cal S}_{1}$ (a barrier) connecting terminals $L$ and $R$ to which a hot scanning tip is approached. This is the configuration of Fig. 3(a) with ${\cal S}_{2}=\mathbb{1}$ and ${\cal S}_{1}$ depending on a constant transmission probability. The kinetic phase accumulated in multiple reflections between the barrier and the tip (separated by a distance $x$) introduces an energy dependent $2k(E)|x|$-periodic interference pattern in the three-terminal scattering matrix (composed of the barrier and the tip) sufficient to break the electron-hole symmetry [242]. Meanwhile, the inversion symmetry is broken geometrically by the position of the tip with respect to the barrier. The interference pattern can be further tuned by modifying the position of the tip. This results in a tunable thermoelectric effect as shown in Fig. 3(b), where the current, $I=I_{L}=-I_{R}$, exhibits a series of oscillations with $x$. A similar non-local oscillating behavior was found in tight-binding simulations of a quantum point contact acting as the isothermal conductor [162]. The tunability of the transport properties of these systems allow furthermore for highly efficient thermoelectric rectification [162, 163]. Unfortunately, the heat to work conversion performance is low in these devices, as a consequence of a considerable fraction of the heat being lost to the terminal without thermoelectric scatterer and the transmission probability being sub-optimal.
@@
The performance can be improved by considering scattering regions with more useful properties ${\cal S}_{\alpha}(E)$ and exploring the capability to engineer the global transmission probabilities of a quantum thermocouple using the electron phase. A system comprised by quantum dots (or double quantum dots) as scattering regions $1$ and $2$, has been proposed [115] as a coherent analog of the inelastic thermocouples [126, 71] discussed in Sec. III.1.1. While this system allows for direct transport between terminals $L$ and $R$, which is in principle detrimental for the thermocouple operation (as discussed in Sec. III), optimizing the different available parameters one finds regimes for which the tip-mediated interference within the conductor becomes destructive and ${\cal T}_{LR}(E)\approx 0$, see Fig. 3(c). Furthermore, this interference can improve the transmission properties of the two junctions giving rise to an enhancement of all performance quantifiers with respect to the inelastic case. In particular, for double quantum dots, the resulting transmissions develop a boxcar-shape [230], attaining efficiencies of $\eta>0.4\eta_{C}$ with high extracted powers (around $70\%$ of the theoretical maximum) and reduced signal to noise ratio [115]. A slight improvement of the performance was also found in tight-binding coherent models of the quantum dot thermocouple configuration just discussed [150], which can again be attributed to suppressed transport between isothermal terminals and the favorable modification of the HLH-L and HRH-R transmissions.
@@
The manipulation of the electron phase via the Aharonov-Bohm effect [245] in three-terminal coherent ring configurations to design quantum thermocouples has also been explored [142, 143, 246, 164]. The transport properties of these devices are easily tunable via the magnetic phase $\phi=2\pi\Phi/\Phi_{0}$ accumulated upon performing a loop around the ring, with $\Phi$ being the magnetic flux piercing the ring and $\Phi_{0}=h/e$ the flux quantum [247, 16]. In these setups the electron-hole symmetry is broken by the kinetic phase accumulated between the junctions connecting the ring to the different terminals [248, 249], with broken time-reversal symmetry introducing the mirror asymmetry (via the non reciprocity of the transmission probabilities ${\cal T}_{ij}(B)\neq{\cal T}_{ji}(B)$) needed for the thermocouple effect, even if the geometric configuration of the system is symmetric and energy-independent [164]. To see this, consider a symmetric ring with all three terminals equally separated by one third of the ring perimeter, so we have ${\cal T}_{\circlearrowright}(\phi)={\cal T}_{12}(\Phi)={\cal T}_{23}(\Phi)={\cal T}_{31}(\Phi)$. A Sommerfeld expansion [250] of the linear response coefficients (valid at low temperatures, $k_{\text{B}}T\ll\mu$ provided the transmission probabilities are smooth around $E=\mu$ [163]), such that $I_{i}\approx 2\sum_{j}[(h^{-1}{\cal T}_{ij}(\mu)(\mu_{i}-\mu_{j})+q_{H}{\cal T}^{\prime}_{ij}(\mu)(T_{i}-T_{j})]$, results in a zero-bias current:
$I_{1}=2q_{H}\frac{{\cal T}^{\prime}_{\circlearrowright}(\mu,\phi){\cal T}_{\circlearrowright}(\mu,{-}\phi){-}{\cal T}^{\prime}_{\circlearrowright}(\mu,{-}\phi){\cal T}_{\circlearrowright}(\mu,\phi)}{{\cal T}_{\circlearrowright}(\mu,\phi)+{\cal T}_{\circlearrowright}(\mu,-\phi)}\Delta T_{H},$ (15)
@@ III.2.1 Dephasing
The relaxation of electrons within the conductor can also be modeled via probe terminals by imposing the boundary conditions $I_{p}=J_{p}=0$: the probe adapts its temperature and electrochemical potential, thus behaving as a thermometer, and enables the definition of an effective local temperature in different regions of the conductor [252, 253, 254, 255, 256, 257], which also shows an interference pattern [163]. The measurement of local dissipation in one dimensional conductors via scanning probe tips has also been described with more involved models combining the scattering approach with a nonlinear Boltzmann equation for the electrons [258].
III.3 Edge states
@@
A way to control which electrons get in contact with the heat source is to use the chirality of quantum Hall edge states [259, 260] in the presence of strong magnetic fields. The quantum Hall regime is indeed a multiterminal effect, and thermoelectric measurements have elegantly demonstrated the chiral propagation of electrons via the energy dissipated in the terminals along the way from the heat source [76, 77, 78]. Measurements of the temperature of a small (a few $\mathrm{\mu m^{2}}$) metallic island coupled via a selected number of edge channels to a dissipative current evidenced the quantization of the thermal conductance [261, 262] and the occurrence of heat Coulomb blockade if the island $RC$ time does not allow to stabilize the charge fluctuations [263, 264].
In addition to the separation of heat and particle flows discussed in the previous sections, the quantum Hall regime achieves the separation of left moving and right moving electrons on opposite sides of the conductor, making it possible to inject heat only on electrons moving in a particular direction [265, 153, 154, 155, 266], see Fig. 4(a). In the absence of interedge scattering, heat flows only along the edge connected to the hot terminal. Along the other edge, the particle current vanishes. In order to generate a finite thermoelectric response, the two edges need to be coupled by an energy-dependent scattering region, see Fig. 4(b). This can be a gate tunable junction like, e.g., a quantum point contact or a quantum dot, or induced by the difference of kinetic phases in a Mach-Zehnder interferometer [156]. Interestingly, if the scatterer is placed only between the heat source and one of the conductor terminals, as shown in Fig. 4(b), the Seebeck and Peltier effects (which are linked by Onsager reciprocity relations in the absence of a magnetic field) are also separated, i.e., one is finite with the other one vanishing [154]: $\Delta T$ generates a current between $L$ and $R$ (hot particles from $H$ are filtered at the scatterer), but a difference in $\mu_{L}-\mu_{R}$ does not generate a thermoelectric heat current in terminal $H$ (particles injected from $L$ never reach $H$). This effect was related to a possible enhancement of the efficiency at maximum power [180], however it was later shown that the balance established by the not scattered channel restores the usual bounds [230, 161].
The same discussion can be applied to helical states in topological insulators, which preserve time-reversal symmetry but separate electrons with opposite spin polarization. In that case, the injected heat is converted into a pure spin current [154, 157, 158].
@@ III.4 Nonequilibrium states
Proposals to observe such effect include the use of quantum Hall edge states that facilitate the preparation of a nonequilibrium distribution: two terminals at different local states (e.g., $f_{C}(E)$ and $f_{H}(E)$) can be connected by a scatterer which mixes their distribution. Electrons in one of the outgoing channels from this scattering region will have a distribution $f_{neq}(E)={\cal T}_{C}(E)f_{C}(E)+{\cal T}_{H}(E)f_{H}(E)$, where ${\cal T}_{i}$ is the transmission probability for electrons injected from terminal $i$ [159, 160]. Electrons in the edge state with the resulting distribution can then be injected into the conductor inducing a finite power if they are scattered at a junction with a thermoelectric response, similarly as the discussion in Sec. III.3, see Fig. 4(c). In this model, the fulfillment of the second law can be related to the dissipative currents within the demon, $J_{H}=-J_{C}$, with $\dot{S}_{d}=\dot{S}_{C}+\dot{S}_{H}=J_{H}(T_{C}^{-1}-T_{H}^{-1})\geq-\dot{S}_{s}$. The nonequilibrium situation can also be generated by time-dependent pulses [270]. The addition of channels coupled to different phonon baths have also been discussed to have a similar effect that however cannot be related to a single nonequilibrium state [173].
In experimental configurations, the nonequilibrium situation can be a result of different processes acting on different time scales. This is precisely the regime of operation of hot-carrier solar cells, which can be modeled as a multiterminal conductor [203, 271]. Nonthermal distributions can also be found in interacting edge channels [272, 273, 274, 275] and have been shown to improve the thermoelectric efficiency [105]. Mesoscopic conductors are then of help for understanding how to extract useful power out of nonequilibrium states before they relax in practical setups.